OpenStudy (anonymous):

Let u= (2,-1), v=(-2,4) and w=(-5,7). Decide s and t so that the vector su+tv is twice as long as vector w but in an oppsite direction.

OpenStudy (anonymous):

I've gotten somewhere (I think). su+tv=(2s-5t;7t-s) But how do we get to know the coordinates of (the twice as long and oppositely directed) w?

OpenStudy (anonymous):

I'm so terrible at vector geometry it's starting to get embarassing. :c

OpenStudy (anonymous):

first find w when it is twice longer but in opposite way so thats just -2w=(10,-14). the norm of this will be two times the norm of w and is in the opposite direction... so we find s and t if su+tv=-2w s(2,-1)+t(-2,4)=(10,-14) so we'll solve the following system of equation: 2s-2t=10 -s+4t=-14 solving this we'll have s=2 t=-3

OpenStudy (anonymous):

@Noliec don't be :D I bet you have your own field of experty in mathematics

OpenStudy (anonymous):

Really, you're just able to -2*w and get a twice as (oppositely directed) long vector? Buhuhu me, thanks!

OpenStudy (anonymous):

x_x

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Embarassed for calling you over here @FoolForMath, sorry!

OpenStudy (anonymous):

yes first we set this to opposite direction so -w, then we know that it must be two times so I multiplied it to 2 to get twice as long as that. so the norm of this vector will be sqrt(269)=2sqrt(74) which is 2 times the length of w which is sqrt(74)

OpenStudy (anonymous):

No problem :) Read here: http://www.britannica.com/EBchecked/topic/624306/vector

OpenStudy (anonymous):

Moreover, my favorable area would be (non-geometry) analysis and general proofs.

OpenStudy (anonymous):

I shall read.